OpenStudy (anonymous):

If f(x) = x^2 and g(x) = 7/x, find f(x) * g(x)

OpenStudy (freckles):

multiply x^2 and 7/x

OpenStudy (freckles):

\[f \cdot g \\ x^2 \cdot \frac{7}{x}\]

OpenStudy (freckles):

you can simplify this a little

OpenStudy (freckles):

but make sure you state any domain restrictions upon x when you do

OpenStudy (freckles):

like we know we can't divide by what number?

OpenStudy (anonymous):

So 7x? Can't divide by zero

OpenStudy (freckles):

right so you would f*g=7x where x cannot be zero

OpenStudy (freckles):

 you are done by the way